# Tree-level 11D supergravity amplitudes from the pure spinor worldline

### v1

@@ 5. Discussions
In order to evaluate higher-point correlation functions, one needs to construct multi-particle vertex operators of different ghost numbers satisfying the consistency relations (4.10) and (4.9). The computation of scattering amplitudes then follows from the contraction relations obeyed by these operators and their lower-rank analogues. As a consistency check, one could compare the resulting amplitudes with those obtained from the perturbiner method applied to an arbitrary number of particles. We plan to investigate this problem further in a future work.
@@
Another possible direction for future work is the construction of a pure spinor supertwistor description of 11D supergravity using our results as a basis, in analogy with the 10D case [[35]35,,[36]36,,[37]37]. This may in turn lead to an ambitwistorial framework that exploits the power of CFT for computing supergravity scattering amplitudes. We intend to study this in the near future.
@@ Appendix A. Integration by parts
where only the zero-modes of the fields appearing in the last equation contribute. Due to BRST invariance, the amplitude does not depend on the specific choice of regulator. Plugging eqn. (A.6) into (A.2), the integration furnishes exactly the rule of the equation $\langle\lambda^{7}\theta^{9}\rangle=1$ [[39]39,,[$40]$.
@@
n where we have used the path integral formula in (A.2), with the integration measure defined as $d\Phi]=d^{32}\theta\,[d\lambda]\,[d\overline{\lambda}]\,[dr]$. In the final equality above, we have used the fact that $\tilde{U}$ commutes with $\cal N$}, as established in eqns. (A.15) and (A.18). An important final observation is that the multi-particle vertex operator $V_{ij}^{(1)}$ admits an analogous integration by parts procedure. Specifically, gauge transformations acting on the single-particle vertex operators induce a corresponding transformation in the pinching operator. Concretely, the transformed pinching operator assumes the form
$\eqalignno{&\tilde{V}^{(1)}_{ij}=-{1\over k_{ij}}\left[\tilde{U}^{(1)}_{(i},\tilde{U}^{(0)}_{j)}\right].&({\hbox{A.}21})}$
@@ B.2. Non-linear supergravity: pinching operator
$\eqalignno{T_{MN}{}^{A}&=\partial_{M}E_{N}{}^{A}+(-)^{MN+1}\partial_{N}E_{M}{}^{A}+(-)^{M(N+B)}E_{N}{}^{B}\Omega_{M\,B}{}^{A}+(-)^{BN+1}E_{M}{}^{B}\Omega_{N\,B}{}^{A},&\cr&&({\hbox{B.}24})}$
@@
$\eqalignno{&2D_{(\alpha}h_{\beta)}{}^{A}+(\gamma^{c})_{\alpha}h_{c}{}^{A}+2\Omega_{(\alpha\,\beta)}{}^{A}+(-)^{B+1}2h_{(\alpha}{}^{B}\Omega_{\beta)\,B}{}^{A}&\cr&\ \ \ \ \ \ \ \ \ \ \ \ \ =(-)^{B+1}h_{\beta}{}^{B}T_{\alpha\,B}{}^{A}+h_{\alpha}{}^{C}T_{C\,\beta}{}^{A},&({\hbox{B.}25})}$
@@
$\eqalignno{\{Q,\phi^{a}\}&=(\lambda\gamma^{a}\phi)+\phi^{b}\Omega_{b}{}^{a}+{1\over 2}(\phi\gamma^{a}\phi),&({\hbox{B.}26})\cr\{Q,\phi^{\alpha}\}&=\Omega^{\alpha}-\phi^{\delta}\Omega_{\delta}{}^{\alpha}+\phi^{c}T_{c\delta}{}^{\alpha}\lambda^{\delta},&({\hbox{B.}27})}$
where $\Omega^{\alpha}=\lambda^{\beta}\lambda^{\delta}\Omega_{\beta\,\delta}{}^{\alpha}$, $\Omega_{a}{}^{b}=\lambda^{\alpha}\Omega_{\alpha\,a}{}^{b}$, and $\phi^{A}=\lambda^{\alpha}h_{\alpha}{}^{A}$. Now, we use (B.26) and (B.27) to write the BRST transformations of the two-particle superfields $\phi^{a}_{12}$ and $\phi^{\alpha}_{12}$ as
$\eqalignno{\{Q,\phi^{a}_{12}\}&=(\lambda\gamma^{a}\phi_{12})+\phi^{b}_{1}\Omega_{2,b}{}^{a}+\phi^{b}_{2}\Omega_{1,b}{}^{a}+(\phi_{1}\gamma^{a}\phi_{2}),&({\hbox{B.}28})\cr\{Q,\phi^{\alpha}_{12}\}&=\Omega_{12}^{\alpha}-\phi_{(1}^{\delta}\Omega_{2),\delta}{}^{\alpha}+\phi_{(1}^{c}T_{2),c\,\delta}{}^{\alpha}\lambda^{\delta}.&({\hbox{B.}29})}$
The superfields $\phi^{a}_{12},\phi^{\alpha}_{12},\Omega^{\alpha}_{12}$ appear explicitly in the expression for the pinching operator (2.38), with the field $\phi^{a}_{12}$ appearing in the result of the 4-point amplitude (3.21). It is possible to find an expression for $\phi_{12}^{a}$ in terms of linearized supergravity fields, satisfying the BRST-invariance (B.28). It is given explicitly by
$\eqalignno{\phi_{12}{}^{a}&=\bigg{[}k_{2}^{c}h_{1,c}{}^{d}\Omega_{2,d}{}^{a}+k_{2}^{c}\Omega_{1,c}{}^{d}h_{2,d}{}^{a}-k_{2}^{c}(\phi_{2}\gamma^{a})_{\delta}h_{1,c}{}^{\delta}-h_{2,c}{}^{\beta}(\gamma^{a})_{\delta\beta}T_{1,c\alpha}{}^{\delta}\lambda^{\alpha}&\cr&\ \ \ \ -k_{2}^{c}\phi_{2}^{d}\Omega_{1,cd}{}^{a}-h_{2,c}{}^{d}R_{1,\alpha cd}{}^{a}\lambda^{\alpha}+(1\leftrightarrow 2)\bigg{]}.&({\hbox{B.}30})}$
Indeed, applying the BRST operator on the rhs of (B.30), using the equation of motions (B.19)-(B.16) and (B.7)-(B.12), we see that it satisfies (B.28).
@@
$\eqalignno{&V^{(1)}_{23}={2\over s_{23}}\left(P_{a}\tilde{\phi}^{a}_{23}+d_{\alpha}\tilde{\phi}_{23}^{\alpha}-{1\over 2}\tilde{\phi}_{23,ab}N^{ab}\right).&({\hbox{B.}31})}$
@@
$\eqalignno{\{U_{2}^{(1)},U_{3}^{(1)}\}&=P_{a}\tilde{\chi}_{23}^{a}+d_{\alpha}\tilde{\chi}_{23}^{\alpha}-w_{\alpha}\tilde{\omega}_{23}^{\alpha},&({\hbox{B.}33})}$
@@
$\eqalignno{&\tilde{\chi}_{23}^{a}=\{Q,\hat{\phi}^{a}_{23}\}-(\lambda\gamma^{a}\hat{\phi}_{23}),\ \ \tilde{\chi}_{23}^{\alpha}=[Q,\hat{\phi}^{\alpha}_{23}]-\hat{\Omega}^{\alpha}_{23},\ \ \tilde{\omega}_{23}{}^{\alpha}=\{Q,\hat{\Omega}_{23}^{\alpha}\},&({\hbox{B.}34})}$
@@
$\eqalignno{&\hat{\phi}_{23}^{a}=\phi_{23}^{a}-k_{23}\left(\phi_{2}^{c}h_{3,c}{}^{a}+\phi_{3}^{c}h_{2,c}{}^{a}+\phi_{2}^{\delta}h_{3,\delta}{}^{a}+\phi_{3}^{\delta}h_{2,\delta}{}^{a}\right),&({\hbox{B.}35})\cr&\hat{\phi}_{23}^{\alpha}=\phi_{23}^{\alpha}-k_{23}\left(\phi_{2}^{c}h_{3,c}{}^{\alpha}+\phi_{3}^{c}h_{2,c}{}^{\alpha}+\phi_{2}^{\delta}h_{3,\delta}{}^{\alpha}+\phi_{3}^{\delta}h_{2,\delta}{}^{\alpha}\right),&({\hbox{B.}36})\cr&\hat{\Omega}_{23}{}^{\alpha}=\Omega_{23}{}^{\alpha}-k_{23}\left(\phi^{a}_{(1}\Omega_{2),a}{}^{\alpha}+\phi^{c}_{(1}\lambda^{\beta}T_{2),\beta c}{}^{\alpha}\right).&({\hbox{B.}37})}$
Here, the fields $\phi^{a}_{23},\phi^{\alpha}_{23},\Omega^{\alpha}_{23}$ are the two-particles supergravity fields satisfying the relations (B.28) and (B.29). Next, we compute the left-hand side of eqn. (B.32)
$\eqalignno{&\{Q,V^{(1)}_{23}\}=P_{a}\left(\{Q,\tilde{\phi}^{a}_{23}\}-(\lambda\gamma^{a}\tilde{\phi}_{23})\right)+d_{\alpha}\left([Q,\tilde{\phi}^{\alpha}_{23}]-\tilde{\Omega}^{\alpha}_{23}\right)-w_{\alpha}\{Q,\tilde{\Omega}_{23}^{\alpha}\}.&({\hbox{B.}38})}$
@@
$\eqalignno{&\tilde{\phi}_{23}^{a}=\phi_{23}^{a}-k_{23}\Big{(}\phi_{2}^{c}h_{3,c}{}^{a}+\phi_{3}^{c}h_{2,c}{}^{a}+\phi_{2}^{\delta}h_{3,\delta}{}^{a}+\phi_{3}^{\delta}h_{2,\delta}{}^{a}\Big{)},&({\hbox{B.}39})\cr&\tilde{\phi}_{23}^{\alpha}=\phi_{23}^{\alpha}-k_{23}\Big{(}\phi_{2}^{c}h_{3,c}{}^{\alpha}+\phi_{3}^{c}h_{2,c}{}^{\alpha}+\phi_{2}^{\delta}h_{3,\delta}{}^{\alpha}+\phi_{3}^{\delta}h_{2,\delta}{}^{\alpha}\Big{)},&({\hbox{B.}40})\cr&\tilde{\Omega}_{23}{}^{\alpha}=(\lambda\gamma^{ab})^{\alpha}\tilde{\phi}_{23,ab}=\Omega_{23}{}^{\alpha}-k_{23}\Big{(}\phi^{a}_{(1}\Omega_{2),a}{}^{\alpha}-\phi^{c}_{(1}\lambda^{\beta}T_{2),\beta c}{}^{\alpha}\Big{)}.&({\hbox{B.}41})}$
@@
$\eqalignno{&\{Q,\tilde{\phi}^{a}_{23}\}=\lambda\gamma^{a}\tilde{\phi}_{23}-k_{23}\Big{(}\phi_{(2}^{c}\partial_{c}\phi_{3)}^{a}+\phi_{(2}^{\delta}D_{\delta}\phi_{3)}^{a}-\Omega_{(2}^{\alpha}h^{a}_{3),\alpha}\Big{)},&({\hbox{B.}42})\cr&[Q,\tilde{\phi}^{\alpha}_{23}]=(\lambda\gamma^{ab})^{\alpha}\tilde{\phi}_{23,ab}-k_{23}\Big{(}\phi_{(2}^{c}\partial_{c}\phi_{3)}^{\alpha}+\phi_{(2}^{\delta}D_{\delta}\phi_{3)}^{\alpha}-\Omega_{(2}^{\beta}h^{\alpha}_{3),\beta}\Big{)}.&({\hbox{B.}43})}$

### v2

@@ 5. Discussions
To evaluate higher-point correlation functions, one must construct multi-particle vertex operators of various ghost numbers that satisfy the consistency relations (4.10) and (4.11). For example, the vertex operators $U_{23}^{(0)}$ and $V_{234}^{(1)}$ introduced in section 4 can be derived from the (anti)commutation relations obeyed by the one- and two-particle vertex operators appearing on the right-hand side of eqns. (4.10) and (4.11). Their explicit forms follow directly from eqns. (2.25), (2.27), and (2.38). Since the relevant expressions are BRST-exact by construction, a straightforward comparison yields the results for $U_{23}^{(0)}$ and $V_{234}^{(1)}$. Scattering amplitudes may then be extracted from the contraction relations satisfied by these operators and their lower-rank counterparts. As a non-trivial check, the resulting amplitudes can be compared with those obtained from the perturbiner method for an arbitrary number of external particles.
In principle, this procedure extends to the computation of any tree-level pure spinor correlator in 11D. However, its implementation might be obstructed by subtleties associated with the pole $\eta$ in the integrated vertex operator. These subtleties originate from the structure of the pure spinor measure $[dZ]$ in the projector $\langle\cdot\rangle$, which schematically scales as
$\eqalignno{[dZ]&\sim\lambda^{-7}\,d^{23}\lambda\,d^{23}\overline{\lambda}\,d^{23}r&({5.1})}$
(see Appendix A for details). Consequently, the integrand must decay more slowly than $\lambda^{-16}\overline{\lambda}^{-23}$. This restriction limits the number of integrated vertex operators that may appear in a correlation function, since these operators explicitly involve negative powers of $\eta$. To illustrate this point, define $\xi={\overline{\lambda}\theta\over\lambda\overline{\lambda}+r\theta}$ which satisfies $Q\xi=1$. A naive argument would then suggest $\langle M\rangle=\langle Q(\xi M)\rangle=0$, where $M=\lambda^{7}\theta^{9}$ represents the top cohomology of the 11D pure spinor BRST charge. This conclusion, however, is only valid if $\xi M$ diverges more slowly than $\lambda^{-16}\overline{\lambda}^{-22}$. In fact, $\xi M$ contains terms of the form $\lambda^{-16}\overline{\lambda}^{-22}\theta^{32}r^{22}$, invalidating the naive reasoning. As in 10D [35], these difficulties can be circumvented by introducing a regularized version of the integrated vertex operator in which the problematic poles are absent. While a full treatment of this regularization lies beyond the scope of the present manuscript, we intend to return to this question in future work.
Once the pure spinor correlator has been evaluated for a given set of supergravity external states, the corresponding physical scattering amplitudes are obtained by projecting the result to ordinary spacetime using the measure $\langle\lambda^{7}\theta^{9}\rangle=1$. This measure was derived in [31] within the pure spinor superfield framework, however, to the best of our knowledge, it has not yet been used in explicit computations. In forthcoming work, we plan to analyze 11D pure spinor expressions compatible with the $\lambda^{7}\theta^{9}$ structure which, together with the $\theta$-expansions of [25], will allow us to compute the 4-point function presented in eqn. (3.22) in ordinary spacetime.
@@
Another possible direction for future work is the construction of a pure spinor supertwistor description of 11D supergravity using our results as a basis, in analogy with the 10D case [[36]36,,[37]37,,[38]38,,[39]39]. This may in turn lead to an ambitwistorial framework that exploits the power of CFT for computing supergravity scattering amplitudes. We intend to study this in the near future.
@@ Appendix A. Integration by parts
where only the zero-modes of the fields appearing in the last equation contribute. Due to BRST invariance, the amplitude does not depend on the specific choice of regulator. Plugging eqn. (A.6) into (A.2), the integration furnishes exactly the rule of the equation $\langle\lambda^{7}\theta^{9}\rangle=1$ [[41]41,,[$42]$.
@@
where we have used the path integral formula in (A.2), with the integration measure defined as $d\Phi]=d^{32}\theta\,[d\lambda]\,[d\overline{\lambda}]\,[dr]$. In the final equality above, we have used the fact that $\tilde{U}$ commutes with $\cal N$}, as established in eqns. (A.15) and (A.18).
An important final observation is that the multi-particle vertex operator $V_{ij}^{(1)}$ admits an analogous integration by parts procedure. Specifically, gauge transformations acting on the single-particle vertex operators induce a corresponding transformation in the pinching operator. Concretely, the transformed pinching operator assumes the form
$\eqalignno{&\tilde{V}^{(1)}_{ij}=-{2\over s_{ij}}\left[\tilde{U}^{(1)}_{(i},\tilde{U}^{(0)}_{j)}\right].&({\hbox{A.}21})}$
@@ B.2. Non-linear supergravity: pinching operator
$\eqalignno{T_{MN}{}^{A}&=\partial_{M}E_{N}{}^{A}+(-)^{MN+1}\partial_{N}E_{M}{}^{A}-(-)^{M(N+B)}E_{N}{}^{B}\Omega_{M\,B}{}^{A}-(-)^{BN+1}E_{M}{}^{B}\Omega_{N\,B}{}^{A},&\cr&&({\hbox{B.}24})}$
@@
$\eqalignno{&2D_{(\alpha}h_{\beta)}{}^{A}+(\gamma^{c})_{\alpha\beta}h_{c}{}^{A}-2\Omega_{(\alpha\,\beta)}{}^{A}-(-)^{B+1}2h_{(\alpha}{}^{B}\Omega_{\beta)\,B}{}^{A}&\cr&\ \ \ \ \ \ \ \ \ \ \ \ \ =(-)^{B+1}h_{\beta}{}^{B}T_{\alpha\,B}{}^{A}+h_{\alpha}{}^{C}T_{C\,\beta}{}^{A},&({\hbox{B.}25})}$
@@
$\eqalignno{\{Q,{\tenit\Phi}^{a}\}&=(\lambda\gamma^{a}{\tenit\Phi})-{\tenit\Phi}^{b}{\tenit\Omega}_{b}{}^{a}+{1\over 2}({\tenit\Phi}\gamma^{a}{\tenit\Phi}),&({\hbox{B.}26})\cr\{Q,{\tenit\Phi}^{\alpha}\}&={\tenit\Omega}^{\alpha}+{\tenit\Phi}^{\delta}{\tenit\Omega}_{\delta}{}^{\alpha}+{\tenit\Phi}^{c}T_{c\delta}{}^{\alpha}\lambda^{\delta},&({\hbox{B.}27})}$
where ${\tenit\Omega}^{\alpha}=\lambda^{\beta}\lambda^{\delta}\Omega_{\beta\,\delta}{}^{\alpha}$, ${\tenit\Omega}_{a}{}^{b}=\lambda^{\alpha}\Omega_{\alpha\,a}{}^{b}$, and ${\tenit\Phi}^{A}=\lambda^{\alpha}h_{\alpha}{}^{A}$. Now, we use eqns. (B.26) and (B.27) to write the BRST transformations of the two-particle superfields ${\tenit\Phi}^{a}_{12}$ and ${\tenit\Phi}^{\alpha}_{12}$ as
$\eqalignno{\{Q,{\tenit\Phi}^{a}_{12}\}&=(\lambda\gamma^{a}{\tenit\Phi}_{12})-\phi^{b}_{1}\Omega_{2,b}{}^{a}-\phi^{b}_{2}\Omega_{1,b}{}^{a}+(\phi_{1}\gamma^{a}\phi_{2}),&({\hbox{B.}28})\cr\{Q,{\tenit\Phi}^{\alpha}_{12}\}&={\tenit\Omega}_{12}^{\alpha}+2\phi_{(1}^{\delta}\Omega_{2),\delta}{}^{\alpha}+2\phi_{(1}^{c}T_{2),c\,\delta}{}^{\alpha}\lambda^{\delta}.&({\hbox{B.}29})}$
In eqns. (B.28), (B.29) we used that ${\tenit\Phi}_{i}^{A}=\phi_{i}^{A}$, ${\tenit\Omega}_{i}^{A}=\Omega_{i}^{A}$ for single-particle superfields. It is possible to find an expression for ${\tenit\Phi}_{12}^{a}$ in terms of the linearized 11D supergravity fields, satisfying the BRST-invariance (B.28). This solution can be written as ${\tenit\Phi}_{12}^{a}={1\over s_{12}}\phi_{12}^{a}$, with $\phi_{12}^{a}$ defined by
$\eqalignno{\phi_{12}^{a}&=\bigg[-k_{2}^{c}h_{1,c}{}^{d}\Omega_{2,d}{}^{a}-k_{2}^{c}\Omega_{1,c}{}^{d}h_{2,d}{}^{a}+k_{2}^{c}(\phi_{2}\gamma^{a})_{\delta}h_{1,c}{}^{\delta}+h_{2,c}{}^{\beta}(\gamma^{a})_{\delta\beta}T_{1,c\alpha}{}^{\delta}\lambda^{\alpha}&\cr&\ \ \ \ +k_{2}^{c}\phi_{2}^{d}\Omega_{1,cd}{}^{a}+h_{2,c}{}^{d}R_{1,\alpha cd}{}^{a}\lambda^{\alpha}+(1\leftrightarrow 2)\bigg].&({\hbox{B.}30})}$
As a check, one can apply the BRST operator on the rhs of eqn. (B.30), the use of the equation of motion (B.16)-(B.19) and (B.7)-(B.12), allows one to conclude that ${\tenit\Phi}_{12}^{a}$ indeed satisfies (B.28). Similarly, the 2-particle superfields ${\tenit\Phi}_{12}^{\beta}$, ${\tenit\Omega}_{12,c}{}^{d}$ can be expressed in the form ${\tenit\Phi}_{12}^{\beta}={1\over s_{12}}\phi_{12}^{\beta}$, ${\tenit\Omega}_{12,c}{}^{d}={1\over s_{12}}\Omega_{12,c}{}^{d}$, where
$\eqalignno{\phi^{\beta}_{12}&=\bigg[k_{1}^{a}\phi_{1}^{\delta}\Omega_{2,a\,\delta}{}^{\beta}+h_{1,a}{}^{\delta}k_{2}^{a}\Omega_{2,\delta}{}^{\beta}+k_{12}^{a}T_{12,a\alpha}{}^{\beta}\lambda^{\alpha}+k_{1}^{a}\phi_{1}^{b}T_{2,ab}{}^{\beta}+\phi_{1}^{b}k_{2}^{a}T_{2,ab}{}^{\alpha}&\cr&+k_{1}^{a}\phi_{1}^{\delta}T_{2,a\delta}{}^{\beta}+\phi_{1}^{\delta}k_{2}^{a}T_{2,a\delta}{}^{\beta}+h_{1,a}{}^{c}k_{2}^{a}T_{2,c\alpha}{}^{\beta}\lambda^{\alpha}+(1\leftrightarrow 2)\bigg],&({\hbox{B.}31})\cr\Omega_{12,c}{}^{d}&=\bigg[k_{1}^{a}\Omega_{1,c}{}^{b}\Omega_{2,a\,b}{}^{d}-\Omega_{1,a\,c}{}^{b}k_{2}^{a}\Omega_{2,b}{}^{d}+k_{12}^{a}\lambda^{\alpha}R_{12,a\alpha\,c}{}^{d}&\cr&+h_{1,a}{}^{\beta}k_{2}^{a}R_{2,\beta\alpha\,c}{}^{d}\lambda^{\alpha}+h_{1,a}{}^{b}k_{2}^{a}R_{2,b\alpha\,c}{}^{d}\lambda^{\alpha}+k_{1}^{a}\phi_{1}^{\delta}R_{2,a\delta\,c}{}^{d}+k_{1}^{a}\phi_{1}^{b}R_{2,ab\,c}{}^{d}&\cr&+\phi_{1}^{b}k_{2}^{a}R_{2,ab\,c}{}^{d}++\phi_{1}^{\delta}k_{2}^{a}R_{2,a\delta\,c}{}^{d}+(1\leftrightarrow 2)\bigg].&({\hbox{B.}32})}$
The superfields ${\phi}^{a}_{12},\phi^{\alpha}_{12},\Omega^{\alpha}_{12}$ appear explicitly in the expression for the pinching operator (2.38), with the field $\phi^{a}_{12}$ appearing in the result of the 4-point amplitude (3.22).
@@
$\eqalignno{&V^{(1)}_{23}={2\over s_{23}}\left(P_{a}\tilde{\phi}^{a}_{23}+d_{\alpha}\tilde{\phi}_{23}^{\alpha}-{1\over 2}\tilde{\phi}_{23,ab}N^{ab}\right).&({\hbox{B.}33})}$
@@
$\eqalignno{{s_{23}\over 2}\{U_{2}^{(1)},U_{3}^{(1)}\}&=P_{a}\tilde{\chi}_{23}^{a}+d_{\alpha}\tilde{\chi}_{23}^{\alpha}-w_{\alpha}\tilde{\omega}_{23}^{\alpha},&({\hbox{B.}35})}$
@@
$\eqalignno{&\tilde{\chi}_{23}^{a}=\{Q,\hat{\phi}^{a}_{23}\}-(\lambda\gamma^{a}\hat{\phi}_{23}),\ \ \tilde{\chi}_{23}^{\alpha}=[Q,\hat{\phi}^{\alpha}_{23}]-\hat{\Omega}^{\alpha}_{23},\ \ \tilde{\omega}_{23}{}^{\alpha}=\{Q,\hat{\Omega}_{23}^{\alpha}\},&({\hbox{B.}36})}$
@@
$\eqalignno{&\hat{\phi}_{23}^{a}=\phi_{23}^{a}-{s_{23}\over 2}\left(\phi_{2}^{c}h_{3,c}{}^{a}+\phi_{3}^{c}h_{2,c}{}^{a}+\phi_{2}^{\delta}h_{3,\delta}{}^{a}+\phi_{3}^{\delta}h_{2,\delta}{}^{a}\right),&({\hbox{B.}37})\cr&\hat{\phi}_{23}^{\alpha}=\phi_{23}^{\alpha}-{s_{23}\over 2}\left(\phi_{2}^{c}h_{3,c}{}^{\alpha}+\phi_{3}^{c}h_{2,c}{}^{\alpha}+\phi_{2}^{\delta}h_{3,\delta}{}^{\alpha}+\phi_{3}^{\delta}h_{2,\delta}{}^{\alpha}\right),&({\hbox{B.}38})\cr&\hat{\Omega}_{23}{}^{\alpha}=\Omega_{23}{}^{\alpha}-s_{23}\left(\phi^{a}_{(1}\Omega_{2),a}{}^{\alpha}+\phi^{c}_{(1}\lambda^{\beta}T_{2),\beta c}{}^{\alpha}\right).&({\hbox{B.}39})}$
Next, we compute the left-hand side of eqn. (B.34)
$\eqalignno{&{s_{23}\over 2}\{Q,V^{(1)}_{23}\}=P_{a}\left(\{Q,\tilde{\phi}^{a}_{23}\}-(\lambda\gamma^{a}\tilde{\phi}_{23})\right)+d_{\alpha}\left([Q,\tilde{\phi}^{\alpha}_{23}]-\tilde{\Omega}^{\alpha}_{23}\right)-w_{\alpha}\{Q,\tilde{\Omega}_{23}^{\alpha}\}.&({\hbox{B.}40})}$
@@
$\eqalignno{&\tilde{\phi}_{23}^{a}=\phi_{23}^{a}-{s_{23}\over 2}\Big(\phi_{2}^{c}h_{3,c}{}^{a}+\phi_{3}^{c}h_{2,c}{}^{a}+\phi_{2}^{\delta}h_{3,\delta}{}^{a}+\phi_{3}^{\delta}h_{2,\delta}{}^{a}\Big),&({\hbox{B.}41})\cr&\tilde{\phi}_{23}^{\alpha}=\phi_{23}^{\alpha}-{s_{23}\over 2}\Big(\phi_{2}^{c}h_{3,c}{}^{\alpha}+\phi_{3}^{c}h_{2,c}{}^{\alpha}+\phi_{2}^{\delta}h_{3,\delta}{}^{\alpha}+\phi_{3}^{\delta}h_{2,\delta}{}^{\alpha}\Big),&({\hbox{B.}42})\cr&\tilde{\Omega}_{23}{}^{\alpha}=(\lambda\gamma^{ab})^{\alpha}\tilde{\phi}_{23,ab}=\Omega_{23}{}^{\alpha}-s_{23}\Big(\phi^{a}_{(1}\Omega_{2),a}{}^{\alpha}-\phi^{c}_{(1}\lambda^{\beta}T_{2),\beta c}{}^{\alpha}\Big).&({\hbox{B.}43})}$
@@
$\eqalignno{&\{Q,\tilde{\phi}^{a}_{23}\}=\lambda\gamma^{a}\tilde{\phi}_{23}-s_{23}\Big(\phi_{(2}^{c}\partial_{c}\phi_{3)}^{a}+\phi_{(2}^{\delta}D_{\delta}\phi_{3)}^{a}-\Omega_{(2}^{\alpha}h^{a}_{3),\alpha}\Big),&({\hbox{B.}44})\cr&[Q,\tilde{\phi}^{\alpha}_{23}]=(\lambda\gamma^{ab})^{\alpha}\tilde{\phi}_{23,ab}-s_{23}\Big(\phi_{(2}^{c}\partial_{c}\phi_{3)}^{\alpha}+\phi_{(2}^{\delta}D_{\delta}\phi_{3)}^{\alpha}-\Omega_{(2}^{\beta}h^{\alpha}_{3),\beta}\Big).&({\hbox{B.}45})}$